OpenStudy (anonymous):

I keep making a mistake somewhere and have no idea where Find the slope of the function at the given value y=-(-x+2)^ (1/2) at x=-5

OpenStudy (anonymous):

WE HAVE NOT LEARNED DERIVATIVES YET FYI. We will be getting to them in a day or so, but we haven't actually learned them yet. Just the fundamental theorem thing with f(a+h) and whatnot

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

@DangerousJesse @dan815

OpenStudy (anonymous):

http://www.math.brown.edu/utra/tangentline.html

OpenStudy (anonymous):

The answer is -3

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

wrong problem

OpenStudy (anonymous):

It is actually root 7 over 14

OpenStudy (anonymous):

Which is worse

OpenStudy (fateal):

o.e lel goo luck with that x3

OpenStudy (anonymous):

thanks >.>

OpenStudy (anonymous):

If you know anyone that can help, by all means

OpenStudy (fateal):

ummm...will....i think that my bro and ex know how to work these stuff o.e but ik that they wont help sorry :/

OpenStudy (anonymous):

lol thanks anyways

OpenStudy (fateal):

np ^^

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Refer to the plot and solution calculated by Mathematica 9 Home Edition.

OpenStudy (anonymous):

We haven't gotten to derivatives yet. Just tangent and secant lines. Is it even possible with just that?

OpenStudy (anonymous):

It has been so long since I've been in school, that I can't remember either way. Sorry.

OpenStudy (anonymous):

Awww....

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're familiar with this equation right? \[\Large f^{\prime}(x) = \frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

we know f(x) f(x) = -(-x+2)^ (1/2) what is f(x+h) equal to?

OpenStudy (anonymous):

f(x+h)=f(-5+h)= -(-(-5+h)+2)^1/2

jimthompson5910 (jim_thompson5910):

don't plug in x = -5 yet

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

I'll let you know when that part comes

OpenStudy (anonymous):

well it equals -(-(x+h)+2)^1/2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

I'm going to convert the 1/2 power to a square root since x^(1/2) = sqrt(x)

OpenStudy (anonymous):

Yep

jimthompson5910 (jim_thompson5910):

so we have this so far